OpenStudy (anonymous):

graph the line: Y = -2/3 x + 3

OpenStudy (anonymous):

you might be interested in this: http://www.wolframalpha.com/input/?t=crmtb01&f=ob&i=Y%20%3D%20-2%2F3%20x%20%2B%203

OpenStudy (ashleyisakitty):

 sorry for crappy drawing

OpenStudy (anonymous):

How did you get that

OpenStudy (ashleyisakitty):

i have a graphing calculator.

OpenStudy (anonymous):

Hmmmm...

OpenStudy (ashleyisakitty):

haha..

OpenStudy (anonymous):

you find the x and y intercepts by letting y=0 and x=0. once you have the two points, graph it. do you need further help?

OpenStudy (anonymous):

Im not sure..Im trying to follow that y=0 and x=0...

OpenStudy (anonymous):

to graph a line, you'll need two points. the easiest points to get is the x and y intercept. i.e.(0,y) and (x,0). to find them, let y=0 or x=0. like this: let x=0 y=0+3=3, so there is a pt, (0,3)

OpenStudy (anonymous):

but how did you get 5

OpenStudy (anonymous):

good point there. wolfram gives 4.5.

OpenStudy (anonymous):

even a 4.5 - how did you get that

OpenStudy (anonymous):

let y=0. 0=-2/3 x +3 solve for x

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

you're welcome :)